OpenStudy (anonymous):

Solve for x. Express the answer in simplest form. 2/1 = 8x - 2 / 9

jagr2713 (jagr2713):

Hey there :D is there answer choices ?

OpenStudy (anonymous):

Nope.

jagr2713 (jagr2713):

Ok first we simplify 2/1 to get what?

OpenStudy (anonymous):

2

jagr2713 (jagr2713):

Goo we we have\[2=8x-\frac{ 2 }{ 9 }\] left ok now what do we do next?

OpenStudy (anonymous):

Oops, should have specified= \[2 = (8x - 2) / 9\]

jagr2713 (jagr2713):

nope we add 2/9 to both sides and what do we get?

OpenStudy (anonymous):

http://prntscr.com/7chi80 ;)

jagr2713 (jagr2713):

we add on both sides 2+2/9=?

OpenStudy (anonymous):

? I'm not following.

jagr2713 (jagr2713):

we are tying to x by its self so we add on both sides b.c its negative

jagr2713 (jagr2713):

you get it now @hputerpop

OpenStudy (whpalmer4):

So the actual problem is \[\frac{2}{1} = \frac{8x-2}{9}\]

OpenStudy (whpalmer4):

or \[\frac{2}{1} = 8x - \frac{2}{9}\]

OpenStudy (anonymous):

Yep

jagr2713 (jagr2713):

Or i was doing the 2/1=8x one

jagr2713 (jagr2713):

do u understand @hputerpop

OpenStudy (whpalmer4):

Unclear which one you said "yep" to... \[\frac{2}{1} = \frac{8x-2}{9}\]I would cross-multiply: \[2*9 = 1*(8x-2)\]\[18=8x-2\]\[20=8x\]\[x=\]